1. Directions: Determine each of the following areas and show these graphically. Use probability notation in your final answer. NOTE: Use graphing paper to graph.


1. To the left of z = 1.96 5. Below z = -1.96       9. 1.27≥z≥-2.1

2. At most z = 2.58 6. z≥1       10. z=-0.58

3. At least z = 1.96 7. z≤-1.5

4. To the right of z = 0.33 8. -0.92≤z≤1.75




1
Expert's answer
2022-05-31T08:37:37-0400

1.


"P(Z<1.96)=0.975002"


2.


"P(Z\le2.58)=0.995060"


3.


"P(Z\ge1.96)=1-P(Z<1.96)"

"=0.024998"


4.


"P(Z>0.33)=1-P(Z\le0.33)"

"=0.370700"



5.


"P(Z<-1.96)=0.024998"


6.


"P(Z\ge1)=1-P(Z<1)"

"=0.158655"


7.


"P(Z\le-1.5)=0.066807"





8.


"P(-0.92\le Z\le1.75)"

"=P(Z\le1.75)-P(Z<-0.92)"

"=0.9599408-0.1787864"

"=0.781154"


9.


"P(-2.1\le Z\le1.27)"

"=P(Z\le1.27)-P(Z<-2.1)"

"=0.8979577-0.0178644"

"=0.880093"


10.


"P(Z=-0.58)=0"